OpenStudy (anonymous):

How do you convert the angle from radians to degrees? How do you find the complement of the angle? And How do you find the supplement of the angle???

OpenStudy (anonymous):

angle from radians to degrees ----> multiply the angle by 180/pi compliment of angle = 90- angle (compliment of 50 =90-50=40) supplement of angle = 180- angle (supplement of 100=180-100 =80)